WATS the answer for 3x-2y=4 in graphing
Answer by jmorgan22(8)

3x-2y=4

Since 3x does not contain the variable to solve for, move it to the right-hand side of the equation by subtracting 3x from both sides.
-2y=-3x+4

Divide each term in the equation by -2.
-2y/2 = -3x/-2 + 4/-2
Simplify the left-hand side of the equation by canceling the common factors.

Y=-3x/-2 + 4/-2


Simplify the right-hand side of the equation by simplifying each term.
So now you can graph y = 3x/2 – 2 or f(x)= 3x/2 - 2